# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v3

@@ Abstract
The rapid advancements in artificial intelligence (AI), particularly in generative AI and large language models (LLMs), have profoundly impacted the creative industries, enabling more innovative content creation, enhancing workflows, and democratizing access to creative tools. This paper explores these technological shifts, with particular focus on how those that have emerged since our previous review in 2022 have expanded creative opportunities and improved efficiency. These technological advancements have enhanced the capabilities of text-to-image, text-to-video, and multimodal generation technologies. In particular, key breakthroughs in LLMs have established new benchmarks in conversational AI, while advancements in image generators have revolutionized content creation. We also discuss the integration of AI into post-production workflows, which has significantly accelerated and improved traditional processes. Once content has been created, it must be delivered to its audiences the media industry is facing the demands of increased communication traffic due to creative content. We therefore include a discussion of how AI is beginning to transform the way we represent and compress media content. We highlight the trend toward unified AI frameworks capable of addressing and integrating multiple creative tasks, and we underscore the importance of human insight to drive the creative process and oversight to mitigate AI-generated inaccuracies. Finally, we explore AI’s future potential in the creative sector, stressing the need to navigate emerging challenges and to maximize its benefits while addressing the associated risks.
@@ 1 Introduction
The influence of artificial intelligence (AI) has grown dramatically over the past few years, particularly due to the rise of generative AI and large language models (LLMs). These advancements are widely regarded as beneficial by many countries, creating significant opportunities for growth (e.g. as outlined in the UK, by the Authority of the House of Lords [1]). These advances have also had significant direct and indirect impacts on the creative industries, influencing the direction of their growth. Generative AI, for instance, primarily focuses on generating new data that is not identical to the training data yet shares similarities with it. However, the cardinality of the training data can be huge, larger than what any individual human has ever encountered. The resulting output may therefore act as a new source of insipration.
@@
The next breakthrough happened in 2023 when OpenAI unveiled GPT-4, a significantly larger model with estimated 1.8 trillion parameters and improved performance compared to its predecessors [4]. However, this still represents less than 1% of the human brain’s approximately 600 trillion synaptic connections [^3] https://www.rsb.org.uk/biologist-features/ai-versus-the-brain. GPT-4 is a multimodal large language model that can generate responses to both text and images. It incorporates DALL·E 3, enabling it to comprehend a much broader range of nuances and details than earlier versions. In March 2024, Claude 3 Opus by Anthropic was released, boasting multimodal capabilities in generating images, tables, graphs and diagrams. Moreover, Anthropic claims that Claude 3 Opus outperforms GPT-4 in generating human-like dialog and contextually aware responses. These rapid advances have, in turn, led the creative industries to face significant challenges. For example, DMG Media, the Financial Times, and Guardian Media Group have highlighted concerns about the potential impact on print journalism, particularly if AI tools reduce the need for users to click through to news websites, affecting advertising and subscription revenues [1]. There is also concern about ‘AI-generated slop’—low-quality, mass-produced content created by AI that often lacks coherence or originality [^4] https://reutersinstitute.politics.ox.ac.uk/news/ai-generated-slop-quietly-conquering-internet-it-threat-journalism-or-problem-will-fix-itself. It is typically used for spam, search engines, or clickbait, and is criticized for cluttering the internet and undermining genuine human-created content.
@@
The generation of videos is significantly more challenging for AI than generating images. In February 2024, Google announced Gemini 1.5 which had the capability to process approximately 8 times more data than GPT-4, opening opportunities for video and audio processing  [^5] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/. In the same month, OpenAI provided its first preview of Sora, a model capable of generating impressive realistic videos up to 1 minute long. Based on the videos released by OpenAI, Sora appears to outperform other text-to-video models. Sora is currently available to ChatGPT subscribers. A month later, Gemini 1.5 announced its support for native audio understanding in 180$+$ countries. With the emergence of these tools, together with the prospect of further advances, it is clear that video content creation will be a major beneficiary. This will further open up the media landscape for creativity and provide more opportunities for diverse storytellers, while also reducing production time. A recent example is the AI-generated Christmas commercial by Coca-Cola [^6] https://www.youtube.com/watch?v=4RSTupbfGog. Such advertisements overcome the limitations of current technologies by using very short videos with rapid scene transitions, ensuring that any artifacts, such as unnatural fingers, are less apparent.
@@
In this paper, the objective is to reveal to the reader, the latest technology advancements that have emerged since our previous review paper on AI in the creative industries (published in 2022) [8]. Compared to this earlier paper, which was written when most AI technologies were used as support tools, this updated review describes thesignificant disruptive shifts that have emerged over the past 3-4 years driven by generative AI and other recent AI-based technologies. Similar to [8], we first provide a high-level overview of current advanced AI technologies (Section 2), followed by a selection of key creative domain applications (Section 3) where current AI technologies are changing creative practice. Finally, we discuss the challenges and the future potential of AI associated with the creative industries (Section 4).
@@ 2 Current Advanced AI Technologies
This paper provides a review of AI in the creative industries, building on our previous publication in 2022 [8]. The reader is referred to that work for an introduction to AI, basic neurons, convolutional neural networks (CNNs), generative adversarial networks (GANs), recurrent neural networks (RNNs) and deep reinforcement learning (DRL). In this paper, we emphasize four key technologies that have grown in importance since 2022 that have had a significant impact on the creative industries. These are Transformers, Large language models (LLMs), Diffusion Models (DMs), and Implicit Neural Representations (INRs). It is important to note that, while these newer technologies are gaining prominence, those from previous generations remain in widespread use, often in conjunction with the newer ones. For instance, CNNs complement transformers since CNNs effectively capture local features and semantic meaning, while the attention mechanism in transformers capture global dependencies.
One important class within AI that has become dominant since our previous review comprises Foundation models (FMs). These were described by The Stanford Institute for Human-Centered Artificial Intelligence in 2021 [9] as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks”. Foundation models have been enabled by rapid advances in AI-oriented computing power and have been underpinned the emergence and success of Large Language Models, particularly following the launch of ChatGPT by OpenAI in 2022. ChatGPT has become the fastest-growing consumer software application in history [^7] https://www.reuters.com/technology/chatgpt-sets-record-fastest-growing-user-base-analyst-note-2023-02-01/.
@@ 2.1 Transformers
In 2017, Google AI introduced the concept of ‘Transformer’ architectures in their publication ‘Attention Is All You Need’ [10]. This work has since, been instrumental in the development and success of large language models alongside many other applications, including vision understanding [11], and multiple modality learning (e.g., Gato [12]).
Before the advent of transformers, natural language processing (NLP) was performed using recurrent neural networks (RNNs), processing data sequences sequentially. In contrast, the ability of transformers to capture long-range dependencies through self-attention mechanisms that extend across all words in the sequence, meant that the importance of different words could be established globally, understanding relationships regardless of their positions. This context-aware representation enables parallel processing of the entire sequence, making the transformers computationally efficient. A set of several attention layers running in parallel is called Multi-Head Attention.
The Transformer architecture, shown in Fig. 1 (a), comprises Encoder and Decoder sections, similar to many CNN-based generators. However the encoder is now a stack of identical layers, concatenating a multi-head self-attention mechanism and a fully connected feed-forward network. The decoder is also a stack of identical layers, in which each layer has additional sub-layer to perform multi-head attention over the output of the encoder stack.
@@
In 2020, the first successful training of a transformer encoder for image recognition was published [11], reeferred to as a Vision Transformer (ViT). The ViT decomposes an input image into patches, similar to words in a sentence, and processes them through multi-head attention. Additionally, a Multilayer Perceptron (MLP) is employed as the feedforward network. In later work Microsoft introduced a hierarchical division of image inputs and a shifted window approach in their Swin Transformer [16]. This was reported to outperform ViT by 2.4% in ImageNet-22K classification (21,841 different categories). Its version 2 [17] applied a cosine function in the attention module. enabling the scaling of capacity and resolution. More detail on transformer-based object detection is discussed in Section 3.4.2. To date, Swin Transformers have been widely adopted in a range of applications including image restoration [18].
Comprehensive surveys on the use of transformers for image and video processing can be found in [19] and [20], respectively.
@@
In recent years state space models [21, 22], commonly known as ‘Mamba’ have emerged. These are a linear variant of Transformersdistinguished by their linear complexity in attention modeling. They are acknowledged to offer an equivalent or better performance than traditional Transformers, while demanding fewer computational resources and less memory.
@@ 2.2 Large language models
LLMs are based on transformer models using self-attention mechanisms as their core modules. Training LLMs comprises two steps: i) pre-training in an unsupervised learning manner, and ii) fine-tuning to a specific task or prompt-tuning for better user inputs. The models are first ‘pre-trained’ with a large amount of unlabelled text data to learn the meaning of words, and the relationships between those words, before using it to adapt to a downstream task. This is why OpenAI refers to their model as a Generative Pre-trained Transformer (GPT).
Fine-tuning involves training the model on new datasets. The drawback is however that these data need to be large enough to ensure generalization to new tasks. Prompt-tuning and prompt engineering are relatively new disciplines for developing and optimizing prompts to efficiently use language models. Prompts guide the way AI models interpret and respond to user queries. Prompt engineering is the process of structuring text or phrasing that guide the model towards generating the desired output. This relies heavily on trial and error, and an understanding of how the model responds. Prompt-tuning, on the other hand, involves training a small set of parameters before utilizing the LLM, thus requiring a relatively small amount of new data. This approach essentially converts text inputs into task-specific virtual inputs, referred to as tokens, while the pre-trained LLM remains unchanged [25]. The main drawback of prompt-tuning is lack of interpretability. This paradigm has however extended to other domains, such as visual prompt tuning [26]. For a comprehensive survey of LLMs, please refer to [27].
@@ 2.3 Diffusion Models
An important factor in driving the rapid growth of generative AI has been the development of diffusion probabilistic models (referred to as diffusion models (DMs)). The first DM was introduced in 2015 by Sohl-Dickstein et al. [33], using Nonequilibrium Thermodynamics. However, it took a further 5 years for DMs to generate desirable results: the era of DMs began with Denoising Diffusion Probabilistic Models (DDPMs) proposed by Ho et al. [34] in 2020 and Score-based diffusion models proposed by Song et al. [35] in 2021. These involve a simplified process using a denoising autoencoder to approximate Bayesian inference. In brief, the models leverage a diffusion process to learn a probability distribution of the input data. As the name suggests, the data is diffused by gradually adding noise at each iteration step as shown in Fig. 1 (b). A deep neural network (DNN) is then trained to remove this noise, called the denoising process or reverse process. Consequently, the trained model uses random noise to generate data with characteristics similar to those of the training samples. Comparing to GANs, DMs provide higher diversity samples [36] and a training process that is much more stable and does not suffer from mode collapse. DMs are however computationally intensive and require longer training times compared to GANs. The complexity can significantly reduced by training the DMs in latent space. Latent Diffusion Models (LDM) [24] use pretrained networks to convert images to feature maps, and perform training on a low-dimensional space. The diagram of LDM is shown in Fig. 1 (c).
Generating a synthesized sample at random might not be particularly useful, especially for creative industry applications. Therefore, conditional diffusion models have been proposed, supporting a wide range of applications such as text-to-sound, text-to-images, and image-to-videos. For DMs, the conditional distributions are modelled using a conditional denoising autoencoder. Classifier guidance was introduced in [36] to improve the generation of images of a desired class. For example, when we provide the model with information, such as ‘a flower’, the DM will synthesize a variety of flower images, as the word ‘flower’ guides the model toward the latent distribution that is formed by various images of flowers. The work in [37] simply refines the latent space of well-trained unconditional DDPM so that the higher-level semantics of the synthetic samples are similar to the reference (conditioning). The LDM [24] offers more flexible conditional image generators by adding cross-attention layer (referred to Transformers in Section 2.1) to the denoising autoencoder. A survey on the methods and applications of DMs prior to 2024 can be found in [38].
@@ 2.4 Implicit Neural Representations
Implicit Neural Representations (INR), also called neural fields, neural implicits or coordinate-based neural networks, represent input content implicitly through learned functions $F$, as shown in Eq. 3. They can be considered as fields $x$ (represented by a scalar, vector, or a tensor with a value, such as magnetic field in physics) that are fully or partially parameterized by a neural network $\Phi$, typically an MLP [39].
@@
Traditional MLPs employ ReLU (rectified linear unit) for non-linear activation due to its simplicity. However, Sitzmann et al. [41] demonstrated that using periodic functions, such as sinusoids, are more suitable for representing complex natural signals, offering a better fit to the first- and second-order derivatives of the signals. However, this activation can cause ringing artifacts. Saragadam et al. instead proposed using complex Gabor wavelets [42], which learn to represent high frequencies better and simultaneously are robust to noise.
@@
Information Text categorization [83, 84, 85, 86]
Analysis Film analysis [87, 88, 89]

### v4

@@ Abstract
The rapid advancements in artificial intelligence (AI), particularly in generative AI and large language models (LLMs), have profoundly impacted the creative industries, enabling more innovative content creation, enhancing workflows, and democratizing access to creative tools. This paper explores these technological shifts, with particular focus on how those that have emerged since our previous review in 2022 have expanded creative opportunities and improved efficiency. These technological advancements have enhanced the capabilities of text-to-image, text-to-video, and multimodal generation technologies. In particular, key breakthroughs in LLMs have established new benchmarks in conversational AI, while advancements in image generators have revolutionized content creation. We also discuss the integration of AI into post-production workflows, which has significantly accelerated and improved traditional processes. Once content has been created, it must be delivered to its audiences; the media industry is now facing the demands of increased communication traffic due to creative content. We therefore include a discussion of how AI is beginning to transform the way we represent and compress media content. We highlight the trend toward unified AI frameworks capable of addressing and integrating multiple creative tasks, and we underscore the importance of human insight to drive the creative process and oversight to mitigate AI-generated inaccuracies. Finally, we explore AI’s future potential in the creative sector, stressing the need to navigate emerging challenges and to maximize its benefits while addressing the associated risks.
@@ 1 Introduction
The influence of artificial intelligence (AI) has grown dramatically over the past few years, particularly due to the rise of generative AI and large language models (LLMs). These advancements are widely regarded as beneficial by many countries, creating significant opportunities for growth (e.g. as outlined in the UK, by the Authority of the House of Lords [1]). These advances have also had significant direct and indirect impacts on the creative industries, influencing the direction of their growth. Generative AI, for instance, primarily focuses on generating new data that is not identical to the training data yet shares similarities with it. However, the cardinality of the training data can be huge, larger than what any individual human has ever encountered. The resulting output may therefore act as a new source of inspiration.
@@
The next breakthrough happened in 2023 when OpenAI unveiled GPT-4, a significantly larger model with estimated 1.8 trillion parameters and improved performance compared to its predecessors [4]. However, this still represents less than 1% of the human brain’s approximately 600 trillion synaptic connections [^3] https://www.rsb.org.uk/biologist-features/ai-versus-the-brain. GPT-4 is a multimodal large language model that can generate responses to both text and images. It incorporates DALL·E 3, enabling it to comprehend a much broader range of nuances and details than earlier versions. In March 2024, Claude 3 Opus by Anthropic was released, boasting multimodal capabilities in generating images, tables, graphs, and diagrams. Moreover, Anthropic claims that Claude 3 Opus outperforms GPT-4 in generating human-like dialog and contextually aware responses. These rapid advances have, in turn, led the creative industries to face significant challenges. For example, DMG Media, the Financial Times, and Guardian Media Group have highlighted concerns about the potential impact on print journalism, particularly if AI tools reduce the need for users to click through to news websites, affecting advertising and subscription revenues [1]. There is also concern about ‘AI-generated slop’—low-quality, mass-produced content created by AI that often lacks coherence or originality [^4] https://reutersinstitute.politics.ox.ac.uk/news/ai-generated-slop-quietly-conquering-internet-it-threat-journalism-or-problem-will-fix-itself. It is typically used for spam, search engines, or clickbait, and is criticized for cluttering the internet and undermining genuine human-created content.
@@
The generation of videos is significantly more challenging for AI than generating images. In February 2024, Google announced Gemini 1.5, which had the capability to process approximately 8 times more data than GPT-4, opening opportunities for video and audio processing [^5] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/. In the same month, OpenAI provided its first preview of Sora, a model capable of generating impressive realistic videos up to 1 minute long. Based on the videos released by OpenAI, Sora appears to outperform other text-to-video models. Sora is currently available to ChatGPT subscribers. A month later, Gemini 1.5 announced its support for native audio understanding in 180$+$ countries. With the emergence of these tools, together with the prospect of further advances, it is clear that video content creation will be a major beneficiary. This will further open up the media landscape for creativity and provide more opportunities for diverse storytellers, while also reducing production time. A recent example is the AI-generated Christmas commercial by Coca-Cola [^6] https://www.youtube.com/watch?v=4RSTupbfGog. Such advertisements overcome the limitations of current technologies by using very short videos with rapid scene transitions, ensuring that any artifacts, such as unnatural fingers, are less apparent.
@@
In this paper, the objective is to reveal to the reader the latest technology advancements that have emerged since our previous review paper on AI in the creative industries (published in 2022) [8]. Compared to this earlier paper, which was written when most AI technologies were used as support tools, this updated review describes the significant disruptive shifts that have emerged over the past 3-4 years driven by generative AI and other recent AI-based technologies. Similar to [8], we first provide a high-level overview of current advanced AI technologies (Section 2), followed by a selection of key creative domain applications (Section 3) where current AI technologies are changing creative practice. Finally, we discuss the challenges and the future potential of AI associated with the creative industries (Section 4).
@@ 2 Current Advanced AI Technologies
This paper provides a review of AI in the creative industries, building on our previous publication in 2022 [8]. The reader is referred to that work for an introduction to AI, basic neurons, convolutional neural networks (CNNs), generative adversarial networks (GANs), recurrent neural networks (RNNs) and deep reinforcement learning (DRL). In this paper, we emphasize four key technologies that have grown in importance since 2022 that have had a significant impact on the creative industries. These are Transformers, Large language models (LLMs), Diffusion Models (DMs), and Implicit Neural Representations (INRs). It is important to note that, while these newer technologies are gaining prominence, those from previous generations remain in widespread use, often in conjunction with the newer ones. For instance, CNNs complement transformers since CNNs effectively capture local features and semantic meaning, while the attention mechanism in transformers captures global dependencies.
One important class within AI that has become dominant since our previous review comprises Foundation models (FMs). These were described by The Stanford Institute for Human-Centered Artificial Intelligence in 2021 [9] as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks”. Foundation models have been enabled by rapid advances in AI-oriented computing power and have underpinned the emergence and success of Large Language Models, particularly following the launch of ChatGPT by OpenAI in 2022. ChatGPT has become the fastest-growing consumer software application in history [^7] https://www.reuters.com/technology/chatgpt-sets-record-fastest-growing-user-base-analyst-note-2023-02-01/.
@@ 2.1 Transformers
In 2017, Google AI introduced the concept of ‘Transformer’ architectures in their publication ‘Attention Is All You Need’ [10]. This work has since been instrumental in the development and success of large language models alongside many other applications, including vision understanding [11], and multiple modality learning (e.g., Gato [12]).
Before the advent of transformers, natural language processing (NLP) was performed using recurrent neural networks (RNNs), processing data sequences sequentially. In contrast, the ability of transformers to capture long-range dependencies through self-attention mechanisms that extend across all words in the sequence meant that the importance of different words could be established globally, understanding relationships regardless of their positions. This context-aware representation enables parallel processing of the entire sequence, making the transformers computationally efficient. A set of several attention layers running in parallel is called Multi-Head Attention.
The Transformer architecture, shown in Fig. 1 (a), comprises Encoder and Decoder sections, similar to many CNN-based generators. However, the encoder is now a stack of identical layers, concatenating a multi-head self-attention mechanism and a fully connected feed-forward network. The decoder is also a stack of identical layers, in which each layer has an additional sub-layer to perform multi-head attention over the output of the encoder stack.
@@
In 2020, the first successful training of a transformer encoder for image recognition was published [11], referred to as a Vision Transformer (ViT). The ViT decomposes an input image into patches, similar to words in a sentence, and processes them through multi-head attention. Additionally, a Multilayer Perceptron (MLP) is employed as the feedforward network. In later work Microsoft introduced a hierarchical division of image inputs and a shifted window approach in their Swin Transformer [16]. This was reported to outperform ViT by 2.4% in ImageNet-22K classification (21,841 different categories). Its version 2 [17] applied a cosine function in the attention module. enabling the scaling of capacity and resolution. More detail on transformer-based object detection is discussed in Section 3.4.2. To date, Swin Transformers have been widely adopted in a range of applications including image restoration [18]. Comprehensive surveys on the use of transformers for image and video processing can be found in [19] and [20], respectively.
@@
In recent years, state space models [21, 22], commonly known as ‘Mamba’ have emerged. These are a linear variant of Transformers distinguished by their linear complexity in attention modeling. They are acknowledged to offer an equivalent or better performance than traditional Transformers, while demanding fewer computational resources and less memory.
@@ 2.2 Large language models
LLMs are based on transformer models using self-attention mechanisms as their core modules. Training LLMs comprises two steps: i) pre-training in an unsupervised learning manner, and ii) fine-tuning to a specific task or prompt-tuning for better user inputs. The models are first ‘pre-trained’ with a large amount of unlabelled text data to learn the meaning of words, and the relationships between those words, before using them to adapt to a downstream task. This is why OpenAI refers to their model as a Generative Pre-trained Transformer (GPT).
Fine-tuning involves training the model on new datasets. The drawback is, however, that these data need to be large enough to ensure generalization to new tasks. Prompt-tuning and prompt engineering are relatively new disciplines for developing and optimizing prompts to efficiently use language models. Prompts guide the way AI models interpret and respond to user queries. Prompt engineering is the process of structuring text or phrasing that guides the model towards generating the desired output. This relies heavily on trial and error, and an understanding of how the model responds. Prompt-tuning, on the other hand, involves training a small set of parameters before utilizing the LLM, thus requiring a relatively small amount of new data. This approach essentially converts text inputs into task-specific virtual inputs, referred to as tokens, while the pre-trained LLM remains unchanged [25]. The main drawback of prompt-tuning is the lack of interpretability. This paradigm has, however, extended to other domains, such as visual prompt tuning [26]. For a comprehensive survey of LLMs, please refer to [27].
@@ 2.3 Diffusion Models
An important factor in driving the rapid growth of generative AI has been the development of diffusion probabilistic models (referred to as diffusion models (DMs)). The first DM was introduced in 2015 by Sohl-Dickstein et al. [33], using Nonequilibrium Thermodynamics. However, it took a further 5 years for DMs to generate desirable results: the era of DMs began with Denoising Diffusion Probabilistic Models (DDPMs) proposed by Ho et al. [34] in 2020 and Score-based diffusion models proposed by Song et al. [35] in 2021. These involve a simplified process using a denoising autoencoder to approximate Bayesian inference. In brief, the models leverage a diffusion process to learn a probability distribution of the input data. As the name suggests, the data is diffused by gradually adding noise at each iteration step as shown in Fig. 1 (b). A deep neural network (DNN) is then trained to remove this noise, called the denoising process or reverse process. Consequently, the trained model uses random noise to generate data with characteristics similar to those of the training samples. Comparing to GANs, DMs provide higher diversity samples [36] and a training process that is much more stable and does not suffer from mode collapse. DMs are however computationally intensive and require longer training times compared to GANs. The complexity can be significantly reduced by training the DMs in latent space. Latent Diffusion Models (LDM) [24] use pretrained networks to convert images to feature maps, and perform training on a low-dimensional space. The diagram of LDM is shown in Fig. 1 (c).
Generating a synthesized sample at random might not be particularly useful, especially for creative industry applications. Therefore, conditional diffusion models have been proposed, supporting a wide range of applications such as text-to-sound, text-to-images, and image-to-videos. For DMs, the conditional distributions are modeled using a conditional denoising autoencoder. Classifier guidance was introduced in [36] to improve the generation of images of a desired class. For example, when we provide the model with information, such as ‘a flower’, the DM will synthesize a variety of flower images, as the word ‘flower’ guides the model toward the latent distribution that is formed by various images of flowers. The work in [37] simply refines the latent space of well-trained unconditional DDPM so that the higher-level semantics of the synthetic samples are similar to the reference (conditioning). The LDM [24] offers more flexible conditional image generators by adding cross-attention layers (referred to Transformers in Section 2.1) to the denoising autoencoder. A survey on the methods and applications of DMs prior to 2024 can be found in [38].
@@ 2.4 Implicit Neural Representations
Implicit Neural Representations (INR), also called neural fields, neural implicits or coordinate-based neural networks, represent input content implicitly through learned functions $F$, as shown in Eq. 3. They can be considered as fields $x$ (represented by a scalar, vector, or a tensor with a value, such as a magnetic field in physics) that are fully or partially parameterized by a neural network $\Phi$, typically an MLP [39].
@@
Traditional MLPs employ ReLU (rectified linear unit) for non-linear activation due to its simplicity. However, Sitzmann et al. [41] demonstrated that periodic functions, such as sinusoids, are more suitable for representing complex natural signals, offering a better fit to the first- and second-order derivatives of the signals. However, this activation can cause ringing artifacts. Saragadam et al. instead proposed using complex Gabor wavelets [42], which learn to represent high frequencies better and simultaneously are robust to noise.
@@
Information Analysis Text categorization [83, 84, 85, 86]
Film analysis [87, 88, 89]